DIODE LASER SPECTROSCOPY OF N2–D2O COMPLEX IN THE V2 BEND REGION OF D2O

SONG LI, RUI ZHENG, YU ZHU, YU YANG, and \underline{CHUANXI DUAN}, College of Physical Science and Technology, Central China Normal University, Wuhan 430079, China.

The rovibrational spectrum of N2–D2O complex was measured in the v2 bend region of D2O using a rapid-scan diode laser spectrometer to probe a pulsed supersonic slit jet. Three subbands, namely Ka = 1–0, 0–0, and 0–1, were recorded. All the transitions in the spectrum are doubled, due to D2O tunneling within the complex. The spectrum was analyzed in terms of a standard asymmetric rotor Hamiltonian, yielding band origins, rotational and centrifugal distortion constants for the upper and lower tunneling levels. The band origin of N2–D2O is slightly blue-shfited from the monomer, indicating that the intermolecular bond becomes weaker upon vibrational exciation of the v2 bend mode of D2O. The tunneling dynamics within the complex will be discussed.